# GeoEdit: Geometric Knowledge Editing for Large Language Models

### v1

@@ 5.1 Experimental Results
The overall results are presented in Table 1. Our GeoEdit method yields promising results in LoRA and full fine-tuning, surpassing other baselines. Surprisingly, ROME maintains Reliability and Generality almost unchanged from the original model across both datasets while achieving high Locality (greater than 90). Since the injection of new knowledge typically impacts Locality, this suggests that ROME performs minimal knowledge updating, likely due to its limited parameter edits. Although F-Learning achieves high Reliability and Generality, its Locality performance drops significantly due to the lack of constraints during the forgetting phase, adversely harms the model’s generalization ability. In contrast, our GeoEdit method, which updates parameters based on the direction of new and old knowledge, improves Locality by 7.4% over F-Learning by masking orthogonal updates. Additionally, by classifying different knowledge editing strategies for new-knowledge-related neurons, our method further improves Reliability and Generality. These results demonstrate that GeoEdit achieves a better balance between updating new knowledge and preserving general knowledge.
@@ 5.2 Ablation Study
In GeoEdit, we categorize knowledge updates into three strategies based on the angle between the directions of knowledge retention and knowledge updating: synergistic, orthogonal, and conflict knowledge editing. To evaluate their impact, we disable each strategy and replace it with vanilla fine-tuning on new knowledge. For example, “- Orthogonal” means setting $h_{edit}=h_{new}$ instead of $h_{edit}=0$.
As shown in Table 2, removing any of the editing strategies leads to performance degradation. Notably, excluding the orthogonal knowledge editing strategy significantly reduces locality, from 75.7% to 71.7%, highlighting the close relationship between orthogonal parameter updates and the model’s general knowledge. Moreover, removing the conflict knowledge editing most adversely affects the reliability metric, causing a decline from 85.2% to 82.6%. These results emphasize the importance of each editing strategy.
@@
The performance decline in Table 2 illustrates the effectiveness of our importance-guided fusion mechanism. This approach offers two key advantages: First, it provides neuron-level adaptive weights, ensuring greater precision than manually set strategies. Second, it ensures that parameter updates are influenced not only by the task vector’s magnitude but also by each neuron’s importance. The smaller weight “masks” significant parameter changes for less important neurons, minimizing their impact on the model’s generalizability.
@@
We investigate the effect of the latent dimension $h_{latent}$ in the auto-encoder on model performance. As shown in Table 3, larger latent dimensions (greater than 1024) cause performance collapse, suggesting that they still suffer from angular bias, with task vectors tending toward orthogonality. Conversely, smaller latent dimensions (less than 256) result in excessive dimensionality reduction, hindering effective training of the auto-encoder and leading to performance degradation. Our experiments demonstrate that a latent dimension of 512 strikes an optimal balance, capturing sufficient information while avoiding unnecessary complexity, which ensures effective model editing and strong generalization.
@@ 5.3 Visualization
We visualize the distribution of the angle $\phi$ between task vectors before and after dimensionality reduction using the auto-encoder, as shown in Figure 3. In the original high-dimensional space, the angles are predominantly concentrated around 90 degrees, as high-dimensional vectors tend to be nearly orthogonal. To mitigate this issue, we trained an auto-encoder to reduce the dimensionality of the task vectors. After dimensionality reduction, the angle distribution spans the full range from 0 to 180 degrees, revealing various types of conflicts between old and new knowledge. This motivates the development of different editing strategies based on angles, allowing us to distinguish between updates that correspond to learning new knowledge and those that modify general knowledge. Thus, we propose a direction-aware geometric model editing approach that effectively handles distinct knowledge updates to enhance model editing.
@@ 5.4 Editing Time Analysis
To evaluate the efficiency of our method, we present the average time required to edit 1000 samples in Table 4. We observe that fine-tuning-based methods are significantly faster than location-based methods, as the latter rely heavily on the specific positions of neurons and parameters, increasing complexity and editing time. Among fine-tuning approaches, F-Learning, which follows a two-stage process of forgetting before learning, takes approximately twice as long as Full-FT. In contrast, our method enables the parallel acquisition of old and new knowledge, resulting in training times comparable to Full-FT. The additional time primarily comes from training the auto-encoder (about 2.5 minutes) and editing the task vector (around 68 seconds). Thus, our method requires less time than F-Learning, delivering substantial performance improvements. This advantage becomes more pronounced as the number of samples to edit increases.
@@ Appendix A Datasets and Examples
It’s important to note that old knowledge represents correct real-world facts, while new knowledge is deliberately incorrect, ensuring that the original model has not previously learned it. This setup avoids ambiguity in determining whether the new knowledge was already part of the model’s prior knowledge.
@@ Appendix B Baseline Details
Appendix C Implementation Details of Experiments
Here we will introduce more completion details and settings of experiments. First, we used LLAMA2-7B and LLAMA-7B as the base models, and then we trained the base model on the old knowledge for 3 epochs by full fine-tuning to simulate an original model that has fully learned old knowledge for our experiments. This makes the forgetting operation more reasonable and effective, and at the same time tries to avoid the problem of being unable to determine whether the new knowledge output by the LLM is learned from the data or commanded by itself as mentioned above. During testing, we use a greedy decoding strategy to ensure the uniqueness of the model’s output. All experiments were conducted on a setup using 4 × A100-80G GPUs.
For full-parameter fine-tuning, all experiments were conducted for 3 epochs with a batch size of 4, gradient accumulation steps set to 4, and a learning rate of 5e-5. Specifically, for LoRA fine-tuning, the hyperparameters were set as follows: $r=8$, $\alpha=32$, dropout = 0.05, with the targeting modules being [q_proj, v_proj]. When we used the Deepspeed, we set 4 processes and zero-stage is 2. The experiment settings for the baseline methods, including F-Learning, FT-c, and ROME, follow those outlined in Ni et al. (2024).
@@ Appendix C Implementation Details of Experiments
Appendix D Sensitivity Analysis for Hyperparameters

### v2

@@ 5.1 Experimental Results
The overall results are presented in Table 1. Firstly, ROME maintains high Reliability and Generality across both datasets while achieving excellent Locality (greater than 90). Since the injection of new knowledge typically impacts Locality, this suggests that ROME performs minimal knowledge updating, likely due to its limited parameter edits. In contrast, F-Learning shows a significant drop in Locality due to the lack of constraints during the forgetting phase, negatively impacting generalization. Our GeoEdit method outperforms fine-tuning-based methods, improving locality by 7.4% over F-Learning. Additionally, by classifying different knowledge editing strategies for new-knowledge-related neurons, our method further improves Reliability and Generality.
AlphaEdit achieves the best performance among locate-and-edit-based methods and outperforms GeoEdit in most cases. This is because AlphaEdit requires the use of an additional 100,000 Wikipedia entries to enhance general knowledge encoding and editing accuracy. In contrast, GeoEdit achieves its performance without relying on any external data. Furthermore, due to the flexibility of fine-tuning methods, we can effectively combine GeoEdit with AlphaEdit (which edits the parameters of the MLP layer, while GeoEdit targets the parameters of the attention layer), creating a complementary approach that further enhances performance.
@@ 5.2 Ablation Study
In GeoEdit, knowledge updates are categorized into synergistic, orthogonal, and conflict editing strategies, based on the angle between knowledge retention and updating directions. To evaluate their impact, we disable each strategy and replace it with vanilla fine-tuning on new knowledge. For example, “- Orthogonal” means setting $h_{edit}=h_{new}$ instead of $h_{edit}=0$. As shown in Table 2, removing any strategy results in performance degradation. Excluding orthogonal editing significantly reduces locality, from 75.7% to 71.7%, while removing conflict editing lowers the reliability metric from 85.2% to 82.6%. These findings underscore the importance of each editing strategy.
@@
The performance decline in Table 2 highlights the effectiveness of our importance-guided fusion. This approach provides two key benefits: it offers neuron-level adaptive weights for greater precision and ensures that parameter updates are influenced by both the task vector’s magnitude and each neuron’s importance. The smaller weights “masks” significant changes for less important neurons, minimizing their impact on the model’s generalizability.
@@
We investigate the effect of the latent dimension $h_{latent}$ in the auto-encoder on model performance. As shown in Table 3, both larger latent dimensions (greater than 1024) and very small latent dimensions (less than 256) lead to performance collapse. This is because the large training loss of the auto-encoder results in poor t-SNE dimensionality reduction, ultimately affecting the accuracy of angular calculations. Our experiments demonstrate that a latent dimension of 512 strikes an optimal balance, effectively removing noise and extracting key features for calculating the angular distribution, which ensures effective model editing and strong generalization.
@@ 5.3 Visualization
We visualize the angle distribution $\phi$ between task vectors using different dimensionality reduction methods, as shown in Figure 3. In high-dimensional space, angles are primarily concentrated around 90 degrees, indicating near orthogonality. Although directly applying t-SNE alleviates this issue to some extent, the angular distribution remains insufficiently dispersed. By first using an AE for denoising and key feature extraction, followed by t-SNE, we achieve a more uniform distribution that spans the full range from 0 to 180 degrees. This allows us to reveal various types of conflicts between old and new knowledge. This motivates the development of editing strategies based on angles, enabling us to distinguish between updates that correspond to learning new knowledge and those that modify general knowledge. The results of the ablation study on different dimensionality reduction methods are provided in Appendix D.1.
@@ 5.4 Editing Time Analysis
Table 4 shows the average time to edit 1000 samples. We find that the editing time of fine-tuning methods is comparable to that of location-based methods, thanks to the use of PEFT techniques and the avoidance of the complex location-based process. Among fine-tuning approaches, F-Learning, which follows a two-stage process of forgetting before learning, takes approximately twice as long as Full-FT. In contrast, our method enables the parallel acquisition of old and new knowledge, resulting in training times comparable to Full-FT. Thus, our method requires less time than F-Learning, delivering substantial performance improvements.
@@ Appendix A Datasets and Examples
It’s important to note that old knowledge represents correct real-world facts, while new knowledge is deliberately incorrect, ensuring that the original model has not previously learned it. This setup avoids ambiguity in determining whether the new knowledge was already part of the model’s prior knowledge Shi et al. (2024).
@@ Appendix B Baseline Details
Appendix C Implementation Details
Here we will introduce more completion details and settings of experiments. First, we used LLAMA2-7B and LLAMA-7B as the base models, and then we trained the base model on the old knowledge for 3 epochs by full fine-tuning to simulate an original model that has fully learned old knowledge for our experiments. This makes the forgetting operation more reasonable and effective, and at the same time tries to avoid the problem of being unable to determine whether the new knowledge output by the LLM is learned from the data or commanded by itself as mentioned above.
We use LoRA to enhance the efficiency of fine-tuning, the hyperparameters were set as follows: $r=8$, $\alpha=32$, dropout = 0.05, with the targeting modules being [q_proj, k_proj, v_proj, o_proj, up_proj, down_proj]. The encoder and decoder consists of 2-layer MLPs with dimensions $[4096\rightarrow 2048,2048\rightarrow 512]$ and $[512\rightarrow 2048,2048\rightarrow 4096]$, respectively, where $d_{latent}$ is set to 512. We set $\lambda$ in Eq. (9) to 0.5, $\phi_{1}$ and $\phi_{2}$ in Eq. (11) to $85^{\circ}$ and $95^{\circ}$, respectively. During testing, we use a greedy decoding strategy to ensure the uniqueness of the model’s output. All experiments were conducted on a setup using 4 × A100-80G GPUs.
@@ Appendix C Implementation Details
Appendix D Additional Results
D.1 Comparison of Different Angle Extraction Methods
Our GeoEdit framework allows using various dimensionality reduction strategies to extract angle information between task vectors. It’s crucial to emphasize that these strategies are simply options or alternatives. The core value of our framework lies in its innovative approach to geometric editing and the proven effectiveness of this method. For example, one could directly apply PCA or t-SNE to the original high-dimensional vectors. However, our empirical results show that the best angle information is achieved by first applying the auto-encoder for denoising, followed by using t-SNE for angle calculation. The related ablation study results are shown in the Table 7.
Method Different Angle Calculation Methods Reliability Generality Locality
F-Learning - 46.9 46.2 72.5
GeoEdit PCA 30.2 27.4 87.7
t-SNE 45.9 46.7 75.5
Auto-Encoder + t-SNE (ours) 46.1 47.2 78.4
Table 7: Comparison of different angle extraction methods.
D.2 Evaluating Fluency and Consistency Scores
In Table 8, we provide the Fluency and Consistency scores for LLAMA-7B and LLAMA2-7B, calculated using the formulas in ROME. Our GeoEdit method consistently outperforms F-Learning in both LoRA-based and full fine-tuning settings, with an average improvement of 33.2 in Fluency and 2.2 in Consistency scores.
Method LLAMA2-7B LLAMA-7B
Fluency ($\uparrow$) Consistency ($\uparrow$) Fluency ($\uparrow$) Consistency ($\uparrow$)
Original model 624.69 26.45 622.94 25.21
LoRA 509.56 18.55 509.01 17.68
Full-FT 251.14 12.33 254.42 11.75
ROME 434.11 8.81 431.86 10.39
RECT 530.82 24.47 532.29 23.32
F-Learning 557.63 26.61 556.02 25.76
AlphaEdit 581.58 30.51 581.68 28.52
GeoEdit 585.98 29.81 584.29 28.90
Table 8: Comparison of different methods for LLAMA2-7B and LLAMA-7B fluency and consistency.
D.3 Results on Different Backbone Models
We have conducted additional experiments using different backbone models, including GPT2-XL (1.5B), Qwen 2.5 (7B), and LLaMA3 (8B). The results in Table 9 show that GeoEdit consistently outperforms RECT and F-Learning across the five key metrics for each backbone model, demonstrating its generalizability across different LLM architectures.
Method Backbone Reliability ($\uparrow$) Generality ($\uparrow$) Locality ($\uparrow$) Fluency ($\uparrow$) Consistency ($\uparrow$)
RECT 63.35 41.55 25.99 529.66 26.67
F-Learning GPT2-XL 64.51 42.56 30.29 544.73 32.34
GeoEdit 66.19 44.43 39.55 575.11 34.92
RECT - 71.80 47.10 29.46 590.31 27.97
F-Learning Qwen 2.5 75.06 48.66 35.31 585.06 27.46
GeoEdit 74.37 50.15 45.41 611.22 30.94
RECT 66.74 43.78 27.38 558.38 26.03
F-Learning LLaMA3 70.69 47.73 33.20 575.16 29.97
GeoEdit 71.42 48.63 41.43 602.38 32.15
Table 9: Comparison of different methods with varying backbones for various metrics.
D.4 The Effect of GeoEdit on Model Generalization
To assess the impact of GeoEdit on generalization, we evaluated mathematical reasoning ability using GSM8K and MATH, as well as broader knowledge retention using MMLU and NLI. The results, summarized in the Table 10, indicate that AlphaEdit, which utilizes additional Wikipedia data, experiences less degradation on MMLU and NLI. Conversely, GeoEdit shows less decline on GSM8K and MATH, demonstrating its effective retention of general knowledge.
Dataset Original Model (LLaMA2-7B) F-Learning AlphaEdit GeoEdit
GSM8K 3.14 0.80 1.55 1.85
MATH 4.32 0.92 2.17 3.01
MMLU 27.74 17.26 21.61 19.72
NLI 67.37 32.60 46.68 42.94
Table 10: Comparison of different methods on various datasets. The best performing method for each dataset is highlighted in bold.
D.5 Analyzing the Importance of Task Vectors in Different MLP Layers for Knowledge Editing
We conducted experiments to analyze the importance of different layers for knowledge editing. For example, in LLaMA-2-7B (which has 32 layers), we separately edited the parameters in the lower (1-11), middle (12-22), and upper (23-32) layers of the MLP using LoRA. The results on the ZsRE dataset are shown in Table 11.
Method Reliability Generality Locality
GeoEdit 54.68 53.45 81.67
GeoEdit (Lower Layer) 49.12 48.46 75.21
GeoEdit (Middle Layer) 47.42 46.18 76.96
GeoEdit (Upper Layer) 46.13 45.31 74.50
Table 11: Comparison of GeoEdit across different layers.
These results indicate that editing only the top layers yields the poorest performance, suggesting that most of the model’s knowledge is stored in the mid-early MLP layers. This is consistent with findings from ROME, and the important parameters are more concentrated in the lower layers, as shown in Figure 4 of the paper.
D.6 Comparison with In-Context-Learning-Based Methods
We evaluated IKE Zheng et al. (2023) on both benchmarks using LLaMA2-7B, with 32 examples and demonstrations selected based on cosine similarity. The results are shown in Table 12.
Method Benchmark Reliability Generality Locality
F-Learning 84.65 81.51 70.92
IKE ZsRE 77.66 76.50 98.30
GeoEdit 85.21 82.43 75.71
F-Learning 69.53 45.56 28.41
IKE Counterfact 60.42 41.71 97.24
GeoEdit 68.34 46.53 37.73
Table 12: Comparison of F-Learning, IKE, and GeoEdit across different benchmarks. The best performance for each metric is highlighted in bold.
The results show that IKE, which does not modify model parameters, minimizes unintended side effects, achieving the highest locality scores, but its reliability and generality are weaker compared to F-Leanring and GeoEdit. While in-context learning methods appear efficient, they face challenges such as the need for large demonstration corpora and sensitivity to factors like demonstration count, selection method, and prompt formatting. These issues can result in performance variability, making them less stable in practice. In contrast, an edited model, whether fine-tuning-based or locate-and-edit, tends to provide more consistent performance and is generally easier to use.
D.7 Sensitivity Analysis for Hyperparameters